Assume the following information for the Milling Department:

Milling Department Units Beginning work in process inventory 300

Units started into production during the period 6,000

Units completed during the period and transferred to the next department 5,750

Ending work in process inventory ?

Using the weighted-average method, how many units were in ending work in process inventory?
Answer #1

Under Process costing there are two cost flow assumptions:-1) FIFO Method and Weighted average method.

First step in the weighted average method:-Preparing Quantity schedule-(Beginning Work in Process+Units started-Units transferred to next department-Ending work in Process)

Beginning Work in Process+Units started=Units transferred out +Ending Work in Process

Example:--Beginning Work in Process=800

                    Units Started=8400

                    Units completed and transferred out=7500

                   Ending Work in process:800+8400-7500=1700

Ans )       Milling Department

            Milling Department Units Beginning work in process inventory 300

           Add: Units started into production during the period 6,000

           Less: Units completed and transferred to the next department 5,750

                       Ending Work in Process=(300+6000-5750)=550 units

Ending Inventory under weighted average method=550 units